# LookAhead: Preventing DeFi Attacks via Unveiling Adversarial Contracts

### v3

@@ 6.4. RQ3: Practicality
From August 14 to August 18, 2024, we deployed the LookAhead system to a cloud server and conducted a live experiment to measure LookAhead’s performance in real-world scenarios. We use a meticulous review process that require us to 1) Wait for $2$ weeks to ensure attacks have been confirmed, 2) Manually review the potential adversarial contracts identified by LookAhead, 3) Validate the detection results by inspecting follow-up transactions involving the adversarial contracts, 4) Calculate the total amount of financial damage caused by valid DeFi attacks, 5) Measure the time gap between the deployment of an adversarial contract and the identification of the contract by LookAhead.
@@
Out of $39$ adversarial contracts flagged by LookAhead, $23$ involved follow-up transactions, and we found $10$ true positives after going through the manual investigation process that were used by real attackers to perform DeFi attacks. These attacks caused over $50,000$ USD in financial losses. In terms of detection efficiency, except only one case where the deployment and attack transaction were confirmed within the same block, LookAhead produced a response for all the attack events before the attacker actually initiated the adversarial transaction. Additionally, after thorough analysis, we observed that almost all the false positives are related to arbitrage, showing that the behaviour exhibited by arbitrage trading resembles some patterns in adversarial activities. For example, the dependency on flashloans, the high number of token ralated function calls, etc.
7. Threats to Validity
@@
It should be noted that contracts may not necessarily need to be verified at the time of deployment, instead, contract creators might choose to complete the verification some time later. Limited by the API responses given by the blockchain explorers, we can only go as far as pinpointing the date of contract verification without knowing the exact time. Thus we cannot determine whether verification was done for a contract at the time of its deployment or not. However, based on our observations, most verified contracts have the same deployment and verification date, and even $6$ attackers verified their adversarial contracts, speculated to have occurred during deployment. Additionally, due to the scarcity of actual adversarial contract samples, the ratio between the two types of samples in our final dataset is approximately $1:40$. Furthermore, benign contracts were selected based on empirical observations and the assumption that adversarial contracts typically have fewer unique interactions with other accounts. However, there is still a chance that we mistakenly treated undisclosed adversarial contracts as benign contracts during training and testing, which could have an impact on the validity of our results.
@@ 8. Related Work
There have been many studies concerning the detection of abnormal contracts, primarily targeting Ponzi schemes and honeypots. They share similarities in methods which could be categorized into code similarity detection (Bartoletti et al., 2020; Sun et al., 2020), machine learning (Chen et al., 2020; Hu et al., 2022; Ji et al., 2023; Zheng et al., 2023), and symbolic execution (Chen et al., 2021). However, to the best of our knowledge, this is the first academic work that focuses on the detection of adversarial contracts across all types of DeFi attacks. Forta (For, 2024) also proposed a method that uses machine learning techniques to identify adversarial contracts, but their approach is limited to bytecode-level analysis. Aside from Forta, the only similar work is BlockWatchDog proposed by Yang et al. (Yang et al., 2024), which uses cross-contract static dataflow analysis technique to detect adversarial contracts in order to identify reentrancy vulnerabilities. However, their focus is solely on reentrancy attacks, which account for only $6.7\%$ of the attacks in our dataset, and their method relies heavily on knowing the victim addresses in advance.

### v4

@@ 6.4. RQ3: Practicality
From August 14 to August 18, 2024, we deployed the LookAhead system to a cloud server and conducted a live experiment to measure LookAhead’s performance in real-world scenarios. We use a meticulous review process that require us to 1) Wait for $2$ weeks (September 1) to ensure attacks have been confirmed, 2) Manually review the potential adversarial contracts identified by LookAhead, 3) Validate the detection results by inspecting follow-up transactions involving the adversarial contracts, 4) Calculate the total amount of financial damage caused by valid DeFi attacks, 5) Measure the time gap between the deployment of an adversarial contract and the identification of the contract by LookAhead.
@@
Out of $39$ adversarial contracts flagged by LookAhead, $23$ involved follow-up transactions, in which we found $10$ true positives after going through the manual investigation process that were used by real attackers to perform DeFi attacks. These attacks caused over $50,000$ USD in financial losses. In terms of detection efficiency, except only one case where the deployment and attack transaction were confirmed within the same block, LookAhead produced a response for all the attack events before the attacker actually initiated the adversarial transaction. Additionally, after thorough analysis, we observed that almost all the false positives are related to arbitrage, showing that the behaviour exhibited by arbitrage trading resembles some patterns in adversarial activities. For example, the dependency on flashloans, the high number of token ralated function calls, etc.
7. Discussion
@@
Threats to Validity. Limited by the API responses given by the blockchain explorers, we can only go as far as pinpointing the date of contract verification without knowing the exact time. Thus we cannot determine whether verification was done at the time of deployment or not. Additionally, due to the scarcity of actual adversarial contract samples, the ratio between the two types of samples in our final dataset is approximately $1:40$. This imbalance could affect the performance of our models. Furthermore, benign contracts were selected based on empirical observations and the assumption that adversarial contracts typically have fewer unique interactions with other accounts. There is still a chance that we mistakenly treated undisclosed adversarial contracts as benign contracts during training and testing, which could have an impact on the validity of our results. Smart contracts can be created through internal transactions triggered during the execution of a normal transaction. Those contracts are not included in our dataset.
Limitations. Our current method has certain limitations in terms of the spectrum of adversarial contracts that LookAhead is capable of detecting. When selecting features, we rely on existing attack patterns seen in DeFi attacks involving financial losses and those in prevalent attack scenarios. Consequently, there is a probability that we might have missed attacks that do not result in financial damages and certain adversarial contracts with novel attack vectors. Furthermore, in the cat-and-mouse game of DeFi attack and defense, attackers may evade our detection by using techniques like code obfuscation to alter the CFG of contract programs or by utilizing KYC-required exchanges as funding sources that makes it harder for our system to recognize adversarial intentions.
@@ 8. Related Work
There have been many studies concerning the detection of abnormal contracts, primarily targeting Ponzi schemes and honeypots. They share similarities in methods which could be categorized into code similarity detection (Bartoletti et al., 2020; Sun et al., 2020), machine learning (Chen et al., 2020; Hu et al., 2022; Ji et al., 2023; Zheng et al., 2023), and symbolic execution (Chen et al., 2021). However, to the best of our knowledge, this is the first academic work that focuses on identifying adversarial contracts, and the first to support detection across all types of DeFi attacks. Forta (For, 2024) also proposed a method that uses machine learning techniques to identify adversarial contracts, but their approach is limited to bytecode-level analysis. Aside from Forta, the only similar work is BlockWatchDog proposed by Yang et al. (Yang et al., 2024), which uses cross-contract static dataflow analysis technique to detect adversarial contracts in order to identify reentrancy vulnerabilities. However, their focus is solely on reentrancy attacks, which account for only $6.7\%$ of the attacks in our dataset, and their method relies heavily on knowing the victim addresses in advance.